# AsEP: Benchmarking Deep Learning Methods for Antibody-specific Epitope Prediction

### v1

@@ A.7 Antibody-antigen complex examples
To compare the epitopes of antibodies in AsEP, we first clustered these complexes by antigen sequences to group together antibodies targeting the same antigen via MMseqs2 Steinegger & Söding (2017). Specifically, we ran MMseqs2 on antigen SEQRES sequences and using the following setup:
@@
To compare epitopes of antibodies sharing the same antigen cluster, we aligned the antigen SEQRES sequences using Clustal Omega Sievers et al. (2011) (download from: http://www.clustal.org/omega/) to obtain a Multiple Sequence Alignment (MSA). Epitopes are mapped to and denoted as the MSA column indices. The epitope similarity between a pair of epitopes is then calculated as the fraction of identical columns. Two epitopes are identified as identical if they share over 0.7 of identical columns.
@@ A.8 Multi-epitope Antigens
A.9 Metrics definition
@@
A.10 Fine-tuning ESMBind on AsEP
The performance reported in the main text for ESMBind is derived by fine-tuning ESM2 on general protein binding sites. We performed a further fine-tuning experiment, fine-tuning it on the presented AsEP dataset and evaluating it on the AsEP test set to enable a more direct comparison of ESMBind to WALLE. Fine-tuning of ESMBind on AsEP was done using the Low-Rank Adaptation method Hu et al. (2021).
@@ A.10 Fine-tuning ESMBind on AsEP
Appendix B Appendix: Link prediction baseline
@@
The optimization process was managed using the early termination functionality provided by the Weights & Biases’ Hyperband method Falkner et al. (2018), with a range of minimum to maximum iterations set from 3 to 27.
The best set of hyperparameters is $2$ GCNConv layers, a batch size of $32$, a weight for positive edges of $54.7$, a weight for the sum of positive links at approximately $75.75e-$ 7, and an edge cutoff of $2.38$. The resulting MCC evaluated on the test set was $0.072$ (standard error: $0.009$) for the bipartite graph link prediction task.
Table S6: Evaluation of WALLE on the bipartite graph link prediction task
Metric Mean (Standard Error)
MCC 0.072 (0.009)
ROC-AUC 0.582 (0.011)
Precision 0.049 (0.008)
Recall 0.167 (0.023)
F1 0.053 (0.007)
@@ C.2 Ablation study: WALLE with simple node encoding
The presented WALLE model utilizes embeddings from large language models, including ESM2 Lin et al. (2022) or IgFoldRuffolo et al. (2023) for representing amino acid sequences, as these models are able to capture the sequential and structural information inherent in protein sequences, providing a rich, context-aware representation of amino acids. To test the effectiveness of such embeddings in this downstream task, we conducted an ablation study where we replaced the embeddings from language models with simple node encodings. Specifically, we evaluated the performance of WALLE when using ‘one-hot’ encoding and ‘BLOSUM62’ encoding for amino acids in both antibody and antigen sequences.
@@ C.3 Ablation study: WALLE with ESM2 embeddings for both antibodies and antigens
We also investigated whether the choice of language models can impact the predictive performance of WALLE; we conducted an ablation study to evaluate the performance of WALLE when both antibodies and antigens are represented using embeddings from the ESM2 language model Lin et al. (2022) while the original model uses AntiBERTy Ruffolo et al. (2023) for antibodies as it is trained exclusively on antibody sequences. This also tests whether a language model trained on general protein sequences can be used for a downstream task like antibody-antigen interaction prediction.
@@
BLOSUM62 Henikoff & Henikoff (1992) encoding involves using the BLOSUM62 matrix, which is a substitution matrix used for sequence alignment of proteins. In this encoding, each residue is represented by its corresponding row in the BLOSUM62 matrix. This method provides a more nuanced representation of residues, reflecting evolutionary relationships and substitution frequencies.
@@ C.4 Hyperparameter tuning
We used the same hyperparameter search space defined in Section A.6 and performed a hyperparameter search as defined in Appendix B for each model variant in the ablation studies. We report the evaluation performance of the tuned model for each variant in Table S7.
@@
Algorithm Encoding MCC AUCROC Precision Recall F1
WALLE Both 0.2097 (0.0195) 0.6351 (0.0126) 0.2346 (0.0183) 0.4217 (0.0279) 0.2580 (0.0178)
WALLE-L Both 0.1593 (0.0155) 0.6124 (0.0109) 0.1750 (0.0109) 0.4696 (0.0243) 0.2371 (0.0137)
WALLE ESM2 0.1955 (0.0212) 0.6219 (0.0137) 0.2280 (0.0188) 0.4103 (0.0291) 0.2553 (0.0188)
WALLE-L ESM2 0.1445 (0.0138) 0.6100 (0.0097) 0.1598 (0.0102) 0.5355 (0.0216) 0.2266 (0.0125)
WALLE One-hot 0.0968 (0.0094) 0.5830 (0.0076) 0.1185 (0.0052) 0.8923 (0.0118) 0.2026 (0.0081)
WALLE BLOSUM 0.0848 (0.010) 0.5739 (0.0081) 0.1182 (0.0055) 0.8401 (0.0151) 0.1993 (0.0083)
@@ Appendix D Impact Statement
Existing methods are compared using uniform metrics and a vast room for improvement can be demonstrated across the board. A novel model is presented, which surpasses its counterparts by an order of magnitude, outperforming them by a factor of ten, validating its graph-based approach utilizing antibody and antigen structural information as well as leveraging protein language models. We provide a foundational framework that invites other researchers to further build upon and refine this baseline model, which we have demonstrated to be a highly effective approach for this task.
@@
print(result_dict) # got {’auc-prc’: tensor(0.5565)}

### v2

@@ A.7 Antibody-antigen complex examples
To compare the epitopes of antibodies in AsEP, we first clustered these complexes by antigen sequences to group together antibodies targeting the same antigen via MMseqs2 (Steinegger & Söding, 2017). Specifically, we ran MMseqs2 on antigen SEQRES sequences and using the following setup:
@@
To compare epitopes of antibodies sharing the same antigen cluster, we aligned the antigen SEQRES sequences using Clustal Omega (Sievers et al., 2011) (download from: http://www.clustal.org/omega/) to obtain a Multiple Sequence Alignment (MSA). Epitopes are mapped to and denoted as the MSA column indices. The epitope similarity between a pair of epitopes is then calculated as the fraction of identical columns. Two epitopes are identified as identical if they share over 0.7 of identical columns.
@@ A.8 Multi-epitope Antigens
A.9 AsEp subset for Benchmarking AlphaFold 2.3 Multimer
As documented in the AlphaFold technical notes (https://github.com/google-deepmind/alphafold/blob/main/docs/technical_note_v2.3.0.md, the AlphaFold version 2.3 Multimer (AF2M) model’s training cutoff date is September 30, 2021. To assess AF2M’s generalizability to novel epitopes not present in the training set, we identified unique epitopes by filtering based on this cutoff date. First, we retrieved all antibody-antigen complex structures from AbDb (snapshot: 2022 September 26) and clustered the antigen sequences using MMseqs2 with the following configuration. The resulting groups are referred to as antigen clusters:
    easy-cluster
    --min-seq-id=0.9  # min sequence identify
    --cov-mode=0      # coverage includes both query (cluster representative)
                      # and target (cluster member) sequence
    --coverage=0.8    # min coverage
Next, we obtained the release date of each AbDb file, cross-referencing with the Protein Data Bank (PDB), and retained only those files in AsEP that were released after the training cutoff date. These files were mapped to their respective antigen clusters.
Finally, we kept only those structures where all cluster members were released after the cutoff date. This filtering process yielded the following 76 structures:
AbDb files:
7pg8_1PΨ7lk4_3PΨ7rd5_1PΨ7tzh_1PΨ7zyi_0PΨ7pnq_0PΨ7rfp_1PΨ7lr4_0PΨ7tuf_1P
7pnm_0P 7rk2_1PΨ7seg_1PΨ7nx3_1PΨ7rk1_1PΨ7amq_0PΨ7sk9_0PΨ7yrf_0PΨ7teq_0P
7usl_0PΨ7vyt_1P 7tpj_0PΨ7ucg_2PΨ7lo7_0PΨ7l7r_1PΨ7vaf_0PΨ7t0r_0PΨ7ttm_0P
7lr3_1PΨ7tuy_0PΨ7a3o_0P 7wk8_0PΨ7ue9_0PΨ7stz_1PΨ7fbi_0PΨ7sfv_0PΨ7tdm_0P
7xw6_0PΨ7fbi_1PΨ7dce_0PΨ7q6c_0P 7s11_2PΨ7rxd_0PΨ7w71_1PΨ7sbg_0PΨ7zwf_0P
7r9d_0PΨ7oh1_0PΨ7po5_0PΨ7qu1_0PΨ7a3p_0P 7ttx_0PΨ7ura_0PΨ7zwm_2PΨ7sjn_0P
7xw7_0PΨ7a3t_0PΨ7tzh_0PΨ7l7r_0PΨ7a0w_0PΨ7u8m_3P 7lo8_0PΨ7zwm_3PΨ7mnl_1P
7n0a_0PΨ7bh8_1PΨ7fci_0PΨ8dke_0PΨ7daa_0PΨ7tug_0PΨ7mrz_0P 7tdn_0PΨ7dkj_0P
7amr_0PΨ7zxf_0PΨ7v61_1PΨ7teb_0P
A.10 AlphaFold2.3 Multimer performance
We used the top one-ranked model generated by AlphaFold2.3 Multimer in the performance calculation in Table S5. See Section A.9 for the list of AbDb files used for this benchmark. The average runtime, including both constructing MSA and predicting structures, is $1.66$ hours (standard deviation: $1.19$ hours, median: $1.27$ hours).
Table S5: AlphaFold2 Multimer performance on novel epitopes after its training cutoff date, 2021 September 30
Method MCC Precision Recall F1
AF2M 0.262 (0.039) 0.276 (0.033) 0.396 (0.043) 0.313 (0.036)
A.11 Metrics definition
@@
A.12 Fine-tuning ESMBind on AsEP
The performance reported in the main text for ESMBind is derived by fine-tuning ESM2 on general protein binding sites. We performed a further fine-tuning experiment, fine-tuning it on the presented AsEP dataset and evaluating it on the AsEP test set to enable a more direct comparison of ESMBind to WALLE. Fine-tuning of ESMBind on AsEP was done using the Low-Rank Adaptation method (Hu et al., 2021).
@@ A.12 Fine-tuning ESMBind on AsEP
Appendix B Appendix-B: Evaluating WALLE with Different Graph Architectures
B.1 Performance of WALLE variants
In this section, we included WALLE’s performance across various graph neural network (GNN) architectures, including GAT, GCN, and GraphSAGE. The results are organized into subsections based on two dataset split strategies: by epitope-to-antigen surface ratio and by epitope group. In Table S7 and Table S8, the protein language models (PLM) used for generating antibody and antigen node embeddings include AntiBERTy, ESM2-35M and ESM2-650M. AntiBERTy is exclusively for generating antibody node embeddings. ESM2-35M and ESM2-650M stand for esm2_t12_35M_UR50D and esm2_t33_650M_UR50D, respectively. The performance values are shown as mean metrics with standard errors provided in parentheses.
Table S7: Performance on epitope prediction
(a) Performance on dataset split by epitope to antigen surface ratio
Method antibody PLM antigen PLM MCC Precision Recall AUCROC F1
GAT AntiBERTy ESM2-35M 0.266 (0.023) 0.287 (0.020) 0.490 (0.027) 0.669 (0.014) 0.325 (0.020)
GCN AntiBERTy ESM2-35M 0.264 (0.021) 0.258 (0.017) 0.534 (0.027) 0.680 (0.014) 0.322 (0.019)
GraphSAGE AntiBERTy ESM2-35M 0.288 (0.022) 0.296 (0.019) 0.521 (0.027) 0.682 (0.014) 0.350 (0.020)
GAT ESM2-35M ESM2-35M 0.254 (0.021) 0.258 (0.016) 0.534 (0.027) 0.669 (0.014) 0.322 (0.018)
GCN ESM2-35M ESM2-35M 0.246 (0.019) 0.234 (0.014) 0.607 (0.025) 0.681 (0.013) 0.310 (0.016)
GraphSAGE ESM2-35M ESM2-35M 0.264 (0.021) 0.269 (0.017) 0.528 (0.027) 0.674 (0.013) 0.327 (0.019)
GAT AntiBERTy ESM2-650M 0.279 (0.024) 0.315 (0.022) 0.437 (0.028) 0.664 (0.014) 0.333 (0.022)
GCN AntiBERTy ESM2-650M 0.305 (0.023) 0.308 (0.019) 0.516 (0.028) 0.695 (0.015) 0.357 (0.021)
GraphSAGE AntiBERTy ESM2-650M 0.294 (0.024) 0.301 (0.020) 0.506 (0.029) 0.684 (0.014) 0.351 (0.022)
GAT ESM2-650M ESM2-650M 0.243 (0.021) 0.291 (0.021) 0.434 (0.028) 0.648 (0.014) 0.297 (0.018)
GCN ESM2-650M ESM2-650M 0.264 (0.022) 0.276 (0.019) 0.509 (0.028) 0.672 (0.014) 0.326 (0.019)
GraphSAGE ESM2-650M ESM2-650M 0.270 (0.021) 0.269 (0.016) 0.533 (0.027) 0.678 (0.013) 0.336 (0.018)
(b) Performance on dataset split by epitope groups
Method antibody PLM antigen PLM MCC Precision Recall AUCROC F1
GAT AntiBERTy ESM2-35M 0.085 (0.017) 0.164 (0.016) 0.225 (0.021) 0.550 (0.010) 0.165 (0.015)
GCN AntiBERTy ESM2-35M 0.122 (0.019) 0.186 (0.018) 0.267 (0.024) 0.576 (0.011) 0.187 (0.017)
GraphSAGE AntiBERTy ESM2-35M 0.113 (0.016) 0.181 (0.015) 0.259 (0.020) 0.568 (0.010) 0.196 (0.015)
GAT ESM2-35M ESM2-35M 0.109 (0.016) 0.175 (0.016) 0.259 (0.023) 0.568 (0.010) 0.177 (0.015)
GCN ESM2-35M ESM2-35M 0.123 (0.017) 0.184 (0.016) 0.345 (0.025) 0.583 (0.011) 0.202 (0.015)
GraphSAGE ESM2-35M ESM2-35M 0.113 (0.017) 0.176 (0.015) 0.279 (0.022) 0.572 (0.011) 0.194 (0.015)
GAT AntiBERTy ESM2-650M 0.125 (0.018) 0.189 (0.017) 0.251 (0.022) 0.575 (0.011) 0.188 (0.016)
GCN AntiBERTy ESM2-650M 0.152 (0.019) 0.207 (0.020) 0.299 (0.025) 0.596 (0.012) 0.204 (0.018)
GraphSAGE AntiBERTy ESM2-650M 0.118 (0.018) 0.206 (0.019) 0.209 (0.021) 0.565 (0.011) 0.176 (0.017)
GAT ESM2-650M ESM2-650M 0.103 (0.017) 0.164 (0.016) 0.240 (0.021) 0.566 (0.011) 0.173 (0.015)
GCN ESM2-650M ESM2-650M 0.138 (0.017) 0.197 (0.018) 0.271 (0.022) 0.585 (0.011) 0.199 (0.016)
GraphSAGE ESM2-650M ESM2-650M 0.111 (0.019) 0.195 (0.019) 0.212 (0.020) 0.559 (0.010) 0.183 (0.017)
Table S8: Performance on bipartite link prediction
(a) Performance on dataset split by epitope to antigen surface ratio
Method antibody PLM antigen PLM MCC Precision Recall AUCROC F1
GAT AntiBERTy ESM2-35M 0.118 (0.010) 0.047 (0.004) 0.393 (0.027) 0.676 (0.014) 0.079 (0.007)
GCN AntiBERTy ESM2-35M 0.114 (0.010) 0.049 (0.004) 0.343 (0.027) 0.657 (0.013) 0.080 (0.007)
GraphSAGE AntiBERTy ESM2-35M 0.122 (0.009) 0.049 (0.004) 0.386 (0.025) 0.675 (0.012) 0.083 (0.006)
GAT ESM2-35M ESM2-35M 0.079 (0.006) 0.020 (0.001) 0.493 (0.027) 0.690 (0.013) 0.037 (0.003)
GCN ESM2-35M ESM2-35M 0.088 (0.008) 0.035 (0.003) 0.301 (0.023) 0.630 (0.011) 0.060 (0.005)
GraphSAGE ESM2-35M ESM2-35M 0.098 (0.008) 0.038 (0.003) 0.333 (0.025) 0.648 (0.012) 0.064 (0.005)
GAT AntiBERTy ESM2-650M 0.121 (0.011) 0.051 (0.005) 0.369 (0.029) 0.670 (0.014) 0.084 (0.008)
GCN AntiBERTy ESM2-650M 0.131 (0.011) 0.062 (0.005) 0.354 (0.029) 0.666 (0.014) 0.095 (0.008)
GraphSAGE AntiBERTy ESM2-650M 0.116 (0.009) 0.044 (0.003) 0.373 (0.027) 0.670 (0.013) 0.077 (0.006)
GAT ESM2-650M ESM2-650M 0.112 (0.009) 0.036 (0.003) 0.484 (0.029) 0.709 (0.014) 0.064 (0.006)
GCN ESM2-650M ESM2-650M 0.119 (0.009) 0.039 (0.003) 0.468 (0.029) 0.707 (0.014) 0.070 (0.006)
GraphSAGE ESM2-650M ESM2-650M 0.112 (0.008) 0.042 (0.003) 0.373 (0.025) 0.668 (0.012) 0.073 (0.005)
(b) Performance on dataset split by epitope groups
Method antibody PLM antigen PLM MCC Precision Recall AUCROC F1
GAT AntiBERTy ESM2-35M 0.034 (0.005) 0.020 (0.003) 0.101 (0.014) 0.542 (0.007) 0.029 (0.004)
GCN AntiBERTy ESM2-35M 0.048 (0.006) 0.020 (0.002) 0.195 (0.022) 0.581 (0.011) 0.034 (0.004)
GraphSAGE AntiBERTy ESM2-35M 0.055 (0.006) 0.026 (0.003) 0.185 (0.019) 0.578 (0.009) 0.042 (0.004)
GAT ESM2-35M ESM2-35M 0.040 (0.005) 0.017 (0.003) 0.222 (0.020) 0.580 (0.009) 0.027 (0.003)
GCN ESM2-35M ESM2-35M 0.044 (0.005) 0.019 (0.003) 0.213 (0.020) 0.578 (0.009) 0.031 (0.004)
GraphSAGE ESM2-35M ESM2-35M 0.047 (0.006) 0.022 (0.003) 0.173 (0.018) 0.570 (0.009) 0.035 (0.004)
GAT AntiBERTy ESM2-650M 0.054 (0.006) 0.022 (0.003) 0.218 (0.022) 0.591 (0.011) 0.037 (0.004)
GCN AntiBERTy ESM2-650M 0.063 (0.007) 0.024 (0.003) 0.265 (0.024) 0.613 (0.012) 0.041 (0.004)
GraphSAGE AntiBERTy ESM2-650M 0.058 (0.007) 0.031 (0.003) 0.174 (0.020) 0.577 (0.010) 0.046 (0.005)
GAT ESM2-650M ESM2-650M 0.044 (0.007) 0.024 (0.004) 0.136 (0.018) 0.558 (0.009) 0.036 (0.005)
GCN ESM2-650M ESM2-650M 0.052 (0.007) 0.024 (0.003) 0.177 (0.020) 0.576 (0.010) 0.039 (0.004)
GraphSAGE ESM2-650M ESM2-650M 0.055 (0.007) 0.028 (0.003) 0.164 (0.018) 0.571 (0.009) 0.043 (0.005)
B.2 Link prediction baseline
@@
The optimization process was managed using the early termination functionality provided by the Weights & Biases’ Hyperband method (Falkner et al., 2018), with a range of minimum to maximum iterations set from 3 to 27.
The best performance on both tasks and dataset splits was achieved using a combination of 2 GCNConv-layer architecture with AntiBERTy and ESM2-650M as the antibody and antigen node embeddings, respectively.
For the epitope prediction task, the best-performing model on the epitope-to-antigen surface ratio dataset split has model parameters that include a weight for positive edges of $140.43$, a weight for the sum of positive links of approximately $1.03\times 10^{-5}$, and an edge cutoff of $3.39$ residues for epitope identification. The decoder applied is an inner-product operator. For the epitope group dataset split, the best-performing model uses an inner-product decoder with a weight for positive edges set to $122.64$, a weight for the sum of positive links of approximately $1.97\times 10^{-7}$, and an edge cutoff of $9.11$ residues for epitope identification. See Table S7 for their performance.
For the bipartite link prediction task, the best-performing model on the epitope-to-antigen surface ratio dataset split shares the same configuration as the best-performing model for the epitope prediction task on this split. On the epitope group dataset split, the best-performing model uses a weight for positive edges of $125.79$, a weight for the sum of positive links of approximately $73.99e-$ 7, and an edge cutoff of $18.56$ residues, also employing an inner-product decoder. See Table S7 for their performance.
@@ C.2 Ablation study: WALLE with simple node encoding
The presented WALLE model utilizes embeddings from large language models, including ESM2 (Lin et al., 2022) or IgFold(Ruffolo et al., 2023) for representing amino acid sequences, as these models are able to capture the sequential and structural information inherent in protein sequences, providing a rich, context-aware representation of amino acids. To test the effectiveness of such embeddings in this downstream task, we conducted an ablation study where we replaced the embeddings from language models with simple node encodings. Specifically, we evaluated the performance of WALLE when using ‘one-hot’ encoding and ‘BLOSUM62’ encoding for amino acids in both antibody and antigen sequences.
@@ C.3 Ablation study: WALLE with ESM2 embeddings for both antibodies and antigens
We also investigated whether the choice of language models can impact the predictive performance of WALLE; we conducted an ablation study to evaluate the performance of WALLE when both antibodies and antigens are represented using embeddings from the ESM2 language model (Lin et al., 2022) while the original model uses AntiBERTy (Ruffolo et al., 2023) for antibodies as it is trained exclusively on antibody sequences. This also tests whether a language model trained on general protein sequences can be used for a downstream task like antibody-antigen interaction prediction.
@@
BLOSUM62 (Henikoff & Henikoff, 1992) encoding involves using the BLOSUM62 matrix, which is a substitution matrix used for sequence alignment of proteins. In this encoding, each residue is represented by its corresponding row in the BLOSUM62 matrix. This method provides a more nuanced representation of residues, reflecting evolutionary relationships and substitution frequencies.
@@ C.4 Hyperparameter tuning
We used the same hyperparameter search space defined in Section A.6 and performed a hyperparameter search as defined in Section B.2 for each model variant in the ablation studies. We report the evaluation performance of the tuned model for each variant in Table S9.
@@
Method Encoding MCC AUCROC Precision Recall F1
WALLE Both 0.264 (0.021) 0.680 (0.014) 0.258 (0.0117) 0.534 (0.027) 0.322 (0.019)
WALLE-L Both 0.159 (0.016) 0.612 (0.011) 0.175 (0.011) 0.470 (0.024) 0.237 (0.014)
WALLE ESM2 0.196 (0.021) 0.622 (0.014) 0.228 (0.019) 0.410 (0.029) 0.255 (0.019)
WALLE-L ESM2 0.145 (0.014) 0.610 (0.010) 0.160 (0.010) 0.536 (0.022) 0.227 (0.013)
WALLE One-hot 0.097 (0.009) 0.583 (0.008) 0.119 (0.005) 0.892 (0.012) 0.203 (0.008)
WALLE BLOSUM 0.085 (0.010) 0.574 (0.008) 0.118 (0.006) 0.840 (0.015) 0.199 (0.008)
@@ Appendix D Impact Statement
Existing methods are compared using uniform metrics and a vast room for improvement can be demonstrated across the board. A novel model is presented, which surpasses its counterparts by an order of magnitude, outperforming them by at least 5 times, validating its graph-based approach utilizing antibody and antigen structural information as well as leveraging protein language models. We provide a foundational framework that invites other researchers to further build upon and refine this baseline model, which we have demonstrated to be a highly effective approach for this task.
@@
print(result_dict)